Find the 43rd term of the sequence an+ 4n-1, where n represents the nth term.
what i think you have to do is an=4(43)-1
But i dont think that is correct and would like your help!
Thank you

Answer by stanbon(75887)

Find the 43rd term of the sequence a(n) = 4n-1, where n represents the nth term.
----
a(43) = 4(43)-1 = 172-1 = 171
========
Cheers,
Stan H.
========